OpenStudy (anonymous):

Which values of x and y will satisfy y = 5x and 12x + 9y = 60?Select one of the options below as your answer: A. x = 1, y = 19 B. x = 35, y = C. x = , y = D. x = , y = E. x = 19, y =

OpenStudy (anonymous):

y = 5x 12x +9y = 60 --> 4x + 3y = 20 sub in 5x for y 4x + 3(5x) = 20 4x + 15x = 20 19x = 20 x = 20/19 y = 5x y = 5*20/19 y = 100/19 this is not one of your choices, is it ?